OpenStudy (anonymous):

simplify ^3square root .027

OpenStudy (dumbcow):

\[\sqrt[3]{.027} = \frac{\sqrt[3]{27}}{\sqrt[3]{1000}} = \frac{3}{10}\]

OpenStudy (anonymous):

is that the root?